### v8

# Movement-Prediction-Adjusted Naïve Forecast

Abstract
This study introduces a movement-prediction-adjusted naïve forecast for time series exhibiting symmetric random walk characteristics, which is applicable after accurate movement predictions are available. Specifically, the original naïve forecast is adjusted by a weighted movement prediction term, where the weights are determined via two parameters derived from the in-sample data: one based on directional accuracy of the movement prediction and the other on the mean absolute increment of the target series. Simulation experiments were conducted across four types of synthetic symmetric random walk series, each with different variance structures. For each time series, diverse movement predictions with predefined directional accuracies were randomly generated, and the resulting forecasts were evaluated via the RMSE, MAE, MAPE, and sMAPE metrics. The results demonstrated a clear monotonic improvement in the forecast performance as the directional accuracy increased. Notably, the adjusted naïve forecast achieved statistically significant improvements even at relatively low directional accuracy levels slightly above 0.50. These findings imply that the movement-prediction-adjusted naïve forecast can serve as an effective second-stage method for forecasting symmetric random walk time series when consistent and accurate movement predictions are provided.
Keywords Naïve forecast  $\cdot$ Movement prediction  $\cdot$ Point forecasting  $\cdot$ Symmetric random walk
1 Introduction
Time series forecasting plays a crucial role in many decision-making processes. The widespread demand for accurate point forecasts across various domains has attracted considerable attention from researchers and practitioners. However, not all time series are equally predictable. Time series exhibiting random walk characteristics pose significant forecasting challenges (Hamilton,, 2020). In particular, financial time series, known for these characteristics and their practical importance in financial markets, have been extensively studied as test cases for forecasting theories and methods. One prominent theory is the efficient market hypothesis (EMH), which suggests that asset prices incorporate all available information, rendering historical prices unpredictable (Fama,, 1970). For a random walk-like series, the naïve forecast, which assumes that the future value at the next time step is equal to the current value, is often regarded as the optimal point forecast.
In support of the EMH, Moosa, (2013) reported that as the volatility of financial time series increases, the root mean square error (RMSE) of the forecasting models increases more quickly than that of the naïve forecast. Moreover, Moosa and Burns, (2014) argued that claims of outperforming the naïve forecast in terms of the RMSE are often misleading, frequently involving introduced dynamics without rigorous statistical validation. Despite numerous attempts to exploit weak structural signals in financial time series via various predictors, existing forecasting methods—including statistical, machine learning, and deep learning approaches—have typically failed to outperform naïve forecasts (Kilian and Taylor,, 2003; Moosa and Burns,, 2016; Thakkar and Chaudhari,, 2021; Petropoulos et al.,, 2022; Ellwanger and Snudden,, 2023; Hewamalage et al.,, 2023; Zeng et al.,, 2023; Beck et al.,, 2025).
From the forecaster’s perspective, the unpredictability of a variable could be attributed to incomplete knowledge of the underlying processes controlling the data, ignorance of the various contributing factors, and imperfections in analytical methods (Laplace,, 2012). In other words, randomness in data is not inherent but rather apparent (Crutchfield and Feldman,, 2003). Although useful information may be contained in the input space, it can be overwhelmed by irrelevant data, making it difficult for current regression methods to exploit such information (John et al.,, 1994; Guyon and Elisseeff,, 2003; Verleysen and François,, 2005). Moreover, identifying the predictors on which the target variable truly depends is a complex task (Granger,, 1969; Stock and Watson,, 2002). However, these theories and empirical evidence reveal only the practical difficulty of finding exploitable patterns; they do not eliminate the possibility of discovering any predictability in the data.
Given the difficulty of generating precise numerical predictions from noisy data in random walk-like time series, researchers have increasingly shifted their focus from point forecasting to movement prediction, which involves predicting whether the value will increase or decrease at the next time step. Consequently, the regression task was reduced to a binary classification task. For a pure random walk, the probability that future movement will increase or decrease is 0.5, and the mutual information between the predictors and the target is zero (Kraskov et al.,, 2004). For a random walk-like time series using abundant predictors, the mutual information can be greater than zero, indicating that there exists some dependency between the predictors and the target (Cover and Thomas,, 2012). Practically, this means that the predictors reduce the uncertainty of the binary outcome. Although this does not guarantee high predictive accuracy, it suggests that, theoretically, one can build a model that performs better than random guessing (Ircio et al.,, 2020). In contrast, for point forecasting in a random walk-like series, the continuous nature of the target means that even if a small gain in mutual information exists, it only marginally reduces the overall uncertainty and does not sufficiently narrow down the precise numerical result (Beraha et al.,, 2019).
Indeed, movement prediction is often considered easier than point forecasting in financial time series (Taylor,, 2008; Moosa,, 2013). Studies in the field of behavioral finance suggest that investor sentiment and psychology can lead to certain patterns (De Bondt and Thaler,, 1985; Lee et al.,, 1991; Baker and Wurgler,, 2007; Barberis et al.,, 1998), which could lead to mutual information between the sentiment predictors and the binary target being greater than zero and making future movements predictable. Therefore, sentiment indicators are often used to improve the directional accuracy in these movement prediction tasks. For example, Weng et al., (2017) incorporated Google news counts and Wikipedia page views alongside historical data to predict the next day’s movement of AAPL stock via support vector machines (SVMs), achieving up to 0.80 directional accuracy. Ma et al., (2023) proposed a multisource aggregated classification (MAC) method for the prediction of stock price movement, incorporating the numerical features and market-driven news sentiments of target stocks, as well as the news sentiments of their related stocks, and the directional accuracy can reach as high as 0.70. Wang et al., (2025) used multimodal data that included news data to train a hybrid model consisting of a long short memory network (LSTM) and transformer for the prediction of stock movement, and the directional accuracy fell within the range of 0.57–0.87. Depending on the complexity of the markets and availability of data, the directional accuracy of financial time series typically ranges between 0.55 and 0.80 (Bustos and Pomares-Quimbaya,, 2020).
When a movement prediction is better than a random guess, it also indicates that the input data contain useful information that has been leveraged effectively to predict future movements. Inspired by the ’less is more’ principle, which suggests that reducing informational complexity can improve decision-making (Gigerenzer and Brighton,, 2009), movement prediction can further serve as a simplified predictor in addition to the naïve forecast to generate the final point forecast, as shown in Figure 1. This indirect approach requires only the naïve forecast and the movement prediction as predictors, eliminating the need for additional noisy predictors. The naïve forecast, which represents the current value, is most closely related to the future value, whereas the movement prediction contributes new directional information not contained within the naïve forecast. Unfortunately, the use of movement prediction alone to improve naïve forecasts remains unexplored in the forecasting community. Previous studies either perform only movement prediction or apply regression directly to noisy data. Although it is often argued that, in many cases, compared with point forecasting, movement prediction is more important for decision-making (Sezer et al.,, 2020), the final decision generally still depends on the estimated value of the target variable. For example, in financial markets, the decision could be to trade at a specific price given a predicted future movement. Thus, even when practitioners rely primarily on predicted movements, they implicitly perform a classification-to-regression conversion by translating the predicted directions into quantitative trading decisions. However, there remains a notable lack of systematic approaches for integrating movement prediction outcomes with naïve forecasts in the context of forecasting random walk-like time series.
Figure 1: Direct and indirect regression approaches
Under these conditions, this study aimed to explore the potential of using movement prediction to enhance point forecasting in random walk-like time series. To this end, we propose a second-stage method that effectively integrates movement prediction with the naïve forecast after accurate and consistent movement predictions have been obtained. The target variable is assumed to follow a symmetric random walk process, which is the basic type of random walk characterized by the absence of drift and equal probabilities of upward and downward movements (Ibe,, 2014). By reformulating the symmetric random walk process to incorporate the future movement of the target variable, the point forecast can naturally be expressed as the original naïve forecast adjusted by the movement prediction term. Hence, it is named the movement-prediction-adjusted naïve forecast. The weight applied to the movement term is determined by two parameters derived from the in-sample data: one based on the directional accuracy of the movement prediction and the other on the mean absolute increment of the target series. The setting of these two parameters involves an inherent trade-off on the prediction performance. While larger parameter values indicate a lower likelihood of achieving improvements through adjustment, these improvements can be substantial; smaller values increase the likelihood of improvement, although such improvements may be marginal or negligible compared with the naïve forecast. To balance this trade-off, the parameters can be empirically estimated by sliding windows on the in-sample data.
Since, in practice, movement prediction can exhibit a wide range of directional accuracies, the adjusted naïve forecast should be evaluated across different accuracy levels to examine the impact of directional accuracy on the performance of the proposed method. Therefore, we use synthetic time series to simulate multiple forecasting scenarios, assuming that for each directional accuracy level, diverse movement predictions are provided. For all the experiments, the evaluation metrics include the RMSE, mean absolute error (MAE), mean absolute percentage error (MAPE), and symmetric mean absolute percentage error (sMAPE), and the original naïve forecast serves as the sole baseline. In this experimental setting, we can evaluate whether the improvement over the baseline by the proposed method is statistically significant.
This study makes three significant contributions to the literature:
  • First, this study offers new insights into the predictability of random walk-like series, presenting a feasible approach to address this longstanding challenge in the field. When meaningful movement prediction is available, the point values of these series, which are traditionally viewed as unpredictable, exhibit conditional predictability.
  • Second, the proposed method bridges movement prediction and point forecasting, which are two distinct tasks in time series forecasting. Although the use of binary classification to guide regression is well established, relying solely on the movement prediction of the target variable for its own point forecasting offers a fresh perspective.
  • Third, this study introduces a clear boundary criterion, which explicitly states that the adjusted naïve forecast outperforms the original naïve forecast when the product of the actual optimal parameter values exceeds half the product of the estimated optimal parameter values. This contribution advances forecasting methodologies by clarifying the conditions required for improved performance.
The remainder of this paper is organized as follows. Section 2 introduces the details of the movement-prediction-adjusted naïve forecast, explains how to derive it and under what conditions it can outperform the original naïve forecast. Section 3 presents a simulation study that assesses the performance of movement-prediction-adjusted naïve forecasts via synthetic datasets in a controlled experimental setting. Section 4 discusses the key findings, implications, and limitations of the study. Finally, Section 5 concludes the paper and outlines future research directions.
2 Method
This section presents the details of the movement-prediction-adjusted naïve forecast. We first derive the general expression of the adjusted naïve forecast based on the definition of symmetric random walks and then explain how to optimize the parameters to improve the forecasting effectiveness and detail the procedure for generating forecasts on the out-of-sample set.
2.1 General Formulation of the Movement-Prediction-Adjusted Naïve Forecast
A symmetric random walk is a stochastic process in which the value at time step $t$ is the sum of the value at the previous time step $t-1$ and the value of an independent random increment ${{\epsilon}_{t}}$, which is generally assumed to follow a normal distribution with a mean of 0 and variance of $\sigma_{t}^{2}$ (De Gooijer et al.,, 2017). The formulation of symmetric random walks is as follows:
${{y}_{t}}={{y}_{t-1}}+{{\epsilon}_{t}},{{\epsilon}_{t}}\sim\mathcal{N}(0,\sigma_{t}^{2}),$ (1)
where ${y}_{t}$ is the point value at time step $t$ and ${y}_{t-1}$ is the point value at time step $t-1$. For random walks exhibiting homoscedasticity, the variance $\sigma_{t}^{2}$ remains constant over time, whereas for those exhibiting heteroscedasticity, the variance $\sigma_{t}^{2}$ varies over time. The symmetric nature of the random increment ${{\epsilon}_{t}}$ ensures that there is no inherent bias in the movement direction of the target variable and that upward and downward movements are equally likely (Ibe,, 2014). Furthermore, the random increment $\epsilon_{t}$ can be expressed as a product of two components: a nonnegative term $|\epsilon_{t}|$ representing the magnitude of the increment and a sign component ${D}_{t}$, which takes a value of $+1$ for upward movement and $-1$ for downward movement, indicating the direction of the increment between two adjacent points. This allows the process to be reformulated as follows:
${{y}_{t}}={{y}_{t-1}}+{{D}_{t}}\cdot|\epsilon_{t}|;|\epsilon_{t}|\geq 0$ (2)
The nonnegative term $|\epsilon_{t}|$ can be further expressed as a product of two terms:
${{y}_{t}}={{y}_{t-1}}+{{D}_{t}}\cdot{\theta_{t}}\cdot\bar{\epsilon};{\theta_{t}}\geq 0,\bar{\epsilon}>0,$ (3)
where $\bar{\epsilon}$ represents the mean absolute increment calculated based on the in-sample data, which can be interpreted as a base magnitude of the increments. The term $\theta_{t}$ is introduced as a nonnegative parameter that controls the magnitude of the increment at time step $t$. Eq. (3) implies that the point forecast of the target variable can be generated if ${{D}_{t}}$ can be predicted and both $\theta_{t}$ and $\bar{\epsilon}$ can be estimated. Consider a simplified scenario in which the magnitude of the increment between two adjacent points is fixed for future time steps; in this case, $\theta_{t}$ can be set to a constant value. Consequently, the point forecast for future time step $t$, denoted as $\hat{y}_{t}$, can be expressed as follows:
${{\hat{y}}_{t}}={{y}_{t-1}}+\hat{D}_{t}\cdot\theta\cdot\bar{\epsilon};\bar{\epsilon}>0,$ (4)
where ${{\hat{D}}_{t}}$ is the movement prediction, $\theta$ is a fixed scalar, and $\theta\in\mathbb{Q}$. Suppose that the movement prediction has been provided, the point forecast generated via Eq. (4) can be regarded as an adjustment of the original naïve forecast through a weighted movement prediction term; hence, it is named the movement-prediction-adjusted naïve forecast.
2.2 Parameter Optimization Using In-Sample Data
According to Eq. (4), the weight of the movement term in the adjusted naïve forecast is the product of two parameters, $\theta$ and $\bar{\epsilon}$, both of which can be optimized based on the in-sample data. Because $\bar{\epsilon}$ is computed directly as the mean absolute increment of the in-sample series, only $\theta$ requires optimization. This optimization is typically achieved by minimizing the mean squared error (MSE) of the adjusted naïve forecast within the in-sample set. Given that the original naïve forecast’s MSE is constant for the in-sample set, minimizing the adjusted naïve forecast’s MSE is equivalent to maximizing their difference:
$\min MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\equiv\max\left(MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\right)$ (5)
Suppose the in-sample set has $N$ time steps, and movement predictions yield $n_{\text{in-sample}}^{\text{correct}}$ correct and $n_{\text{in-sample}}^{\text{incorrect}}$ incorrect directions; then, the directional accuracy of the movement prediction on the in-sample set can be expressed as follows:
$ACC_{\text{in-sample}}=\frac{n_{\text{in-sample}}^{\text{correct}}}{n_{\text{in-sample}}^{\text{incorrect}}+n_{\text{in-sample}}^{\text{correct}}}$ (6)
Accordingly, the MSE difference between two types of naïve forecasts is as follows:
$\Delta MSE_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (7)
$=(4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}})\cdot\bar{\epsilon}_{\text{in-sample}}^{2},$
where $\theta_{\text{in-sample}}$ represents the parameter $\theta$ in the adjusted naïve forecast for the in-sample set and $\bar{\epsilon}_{\text{in-sample}}$ denotes the mean of $|\epsilon_{t}|$ calculated over the in-sample set. A detailed algebraic derivation of the $\Delta MSE_{\text{in-sample}}$ is provided in Appendix A.
According to Eq. (7), the optimization reduces to maximizing the following function concerning parameter $\theta_{\text{in-sample}}$:
$f(\theta_{\text{in-sample}})=4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}},$ (8)
Differentiating Eq. (8) and setting the derivative to zero yields the optimal solution:
Proposition 1 (Optimal parameter solution). The optimal parameter value $\theta_{\text{in-sample}}^{*}$ that maximizes the MSE difference between the adjusted naïve forecast and the original naïve forecast on the in-sample set is given by:
$\theta_{\text{in-sample}}^{*}=2\cdot ACC_{\text{in-sample}}-1$ (9)
This result explicitly reveals that the optimal value of the parameter $\theta$ depends linearly on the directional accuracy of the movement predictions for a sequence with a fixed length. Since the second derivative of $f(\theta_{\text{in-sample}})$ is negative, this optimal solution is a maximum.
Substituting Eq. (9) back into Eq. (6) yields:
Proposition 2 (Performance guarantee). Under the optimal parameter setting using $\theta_{\text{in-sample}}^{*}$, the adjusted naïve forecast achieves a nonnegative improvement over the original naïve forecast in terms of the MSE on the in-sample set:
$\max\Delta MSE_{\text{in-sample}}=\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\geq 0$ (10)
Proposition 2 confirms that as the directional accuracy increases beyond 0.5, the value of $\theta_{\text{in-sample}}^{*}$ increases from 0, leading to progressively greater improvements over the original naïve forecast. At a directional accuracy of 1, the maximum achievable improvement is $\bar{\epsilon}_{\text{in-sample}}^{2}$. Thus, with the optimal parameter setting, the adjusted naïve forecast matches or surpasses the performance of the original naïve forecast within the in-sample dataset:
$MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}\leq MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ (11)
This result underlines the theoretical robustness of the proposed adjustment methodology.
2.3 Uncertainty Analysis and Practical Generation of Out-of-Sample Forecasts
We now examine the performance of the adjusted naïve forecast via out-of-sample data. Suppose that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are used as the parameters of the adjusted naïve forecast on the out-of-sample set, which contains $M$ time steps. Suppose that the movement predictions on this set yield $m_{\text{out-of-sample}}^{\text{correct}}$ correct and $m_{\text{out-of-sample}}^{\text{incorrect}}$ incorrect predictions; then, the directional accuracy of the movement prediction on the out-of-sample set can be expressed as:
$ACC_{\text{out-of-sample}}=\frac{m_{\text{out-of-sample}}^{\text{correct}}}{m_{\text{out-of-sample}}^{\text{incorrect}}+m_{\text{out-of-sample}}^{\text{correct}}}$ (12)
Consequently, the MSE difference between two types of naïve forecasts on the out-of-sample is expressed as:
$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (13)
$=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2},$
where $\bar{\epsilon}_{\text{out-of-sample}}$ denotes the actual mean absolute increment on the out-of-sample set and $\theta_{\text{out-of-sample}}^{*}$ denotes the actual optimal value of $\theta$ for the out-of-sample set, which is derived according to Proposition 1:
$\theta_{\text{out-of-sample}}^{*}=2\cdot ACC_{\text{out-of-sample}}-1.$ (14)
A detailed algebraic expansions of the $\Delta MSE_{\text{out-of-sample}}$ is provided in Appendix B. For the adjusted naïve forecast to outperform the original naïve forecast, the $\Delta MSE_{\text{out-of-sample}}$ should be no less than zero. Therefore, according to Eq. (13), the following inequality must hold:
$\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}\geq\frac{1}{2}\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}$ (15)
Given that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are specific estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, re2spectively, inequality (15) can be generalized as:
Proposition 3 (Out-of-sample performance criterion). The necessary condition for the adjusted naïve forecast to achieve an improvement over the original naïve forecast in terms of the MSE on the out-of-sample set is as follows:
$\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}\geq\frac{1}{2}\cdot\hat{\theta}_{\text{out-of-sample}}\cdot\hat{\bar{\epsilon}}_{\text{out-of-sample}},$ (16)
where $\hat{\theta}_{\text{out-of-sample}}$ represents the estimate of $\theta_{\text{out-of-sample}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ represents the estimate of $\bar{\epsilon}_{\text{out-of-sample}}$.
Proposition 3 establishes a clear boundary criterion when an adjusted naïve forecast can outperform the original naïve forecast, specifically depending on whether the product of the actual optimal parameter values exceeds half the product of the estimated optimal parameter values. We can treat $\theta_{\text{out-of-sample}}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ as random variables, each following a distribution defined over nonnegative values. Accordingly, inequality (16) represents a probabilistic event, the likelihood of which depends on the realized values of these variables. Since $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ cannot be known in advance, the fulfillment of inequality (16) is uncertain. Larger parameter values have a lower likelihood of yielding improvements, although such improvements can be substantial, whereas smaller parameter values increase the likelihood of improvement, albeit potentially marginal or negligible, compared with the naïve forecast. Therefore, this boundary criterion highlights a critical trade-off between the likelihood of achieving improvements and the potential magnitude of these improvements.
In addition to using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, a more conservative approach can be used to generate the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$:
  1. 1.
    Consider a data-generating process in which a sliding window, whose length is equal to that of the out-of-sample set, moves sequentially from the beginning to the end of the in-sample set. Each window can be treated as an "out-of-sample set" in the historical context.
  2. 2.
    Within each window, the directional accuracy is calculated and then translated into the corresponding value of $\theta$ via Eq. (9). As the window moves, a collection of $\theta$ values is generated, denoted as $\Theta_{\text{set}}$, whose bounds can be regarded as the empirical range of $\theta_{\text{out-of-sample}}^{*}$.
  3. 3.
    Similarly, a sliding window of the same length is applied to calculate a set of $\bar{\epsilon}$ values from the in-sample set, denoted as $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, providing an empirical value range of $\bar{\epsilon}_{\text{out-of-sample}}$.
  4. 4.
    Using $\min(\Theta_{\text{set}})$ and $\min(\boldsymbol{\bar{\epsilon}}_{\text{set}})$ as conservative estimates for $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, the adjusted naïve forecast for time step $t$ in the out-of-sample set can be expressed as:
    $\hat{y}_{t}=y_{t-1}+\hat{D}_{t}\cdot\min(\Theta_{\text{set}})\cdot\min(\boldsymbol{\bar{\epsilon}}_{\text{set}})$ (17)
Algorithm 1 describes the process of generating the adjusted naïve forecast on the entire out-of-sample set.
Algorithm 1 Movement-Prediction-Adjusted Naïve Forecast on the Out-of-sample Set
Inputs:
$Y_{\text{in-sample}}$ //    Array of observations in the in-sample set
$Y_{\text{out-of-sample}}$ //    Array of observations in the out-of-sample set
$\hat{D}_{\text{in-sample}}$ //    Movement prediction on the in-sample set
$\hat{D}_{\text{out-of-sample}}$ //    Movement prediction on the out-of-sample set
$N$ //    Number of elements in the in-sample set
$M$ //    Number of elements in the out-of-sample set
Output:
$P$ //    Adjusted naïve forecast on the out-of-sample set
1. Initialize variables:
$D_{\text{in-sample}}\leftarrow[]$ //    Actual movements on the in-sample set
$P\leftarrow[]$
$ACC_{\text{set}}\leftarrow[]$ //    Directional accuracy set
$\Theta_{\text{set}}\leftarrow[]$ //    Theta set
$R_{\text{in-sample}}\leftarrow[]$ //    First-order difference set
$\boldsymbol{\bar{\epsilon}}_{\text{set}}\leftarrow[]$ //    Mean absolute increment set
2. Calculate actual movements:
for $t\leftarrow 1$ to $\text{length}(Y_{\text{in-sample}})$ do
     $\text{diff}\leftarrow Y_{\text{in-sample}}[t]-Y_{\text{in-sample}}[t-1]$ //    First-order difference
     $R_{\text{in-sample}}.\text{append}(\text{diff})$
     if $\text{diff}>0$ then
         $D_{\text{in-sample}}.\text{append}(1)$ //    Upward movement
     else
         $D_{\text{in-sample}}.\text{append}(-1)$ //    Downward or no movement
     end if
end for
3. Calculate directional accuracies:
for $t\leftarrow 1$ to $N-M+1$ do
[…]
     for $i\leftarrow t$ to $t+M-1$ do
         if $D_{\text{in-sample}}[i]==\hat{D}_{\text{in-sample}}[i]$ then
[…]
         end if
     end for
     $ACC_{\text{set}}.\text{append}(\text{Correct\_Predictions}/M)$
end for
4. Compute the movement weights:
for each $\text{each }ACC\in ACC_{\text{set}}$ do
     $\theta\leftarrow 2\cdot\text{ACC}-1$
     $\Theta_{\text{set}}.\text{append}(\theta)$
end for
5. Compute the mean absolute increments:
for $t\leftarrow 1$ to $N-M+1$ do
     $\bar{\epsilon}\leftarrow\text{mean}(|R_{\text{in-sample}}[t:t+M-1]|)$
     $\boldsymbol{\bar{\epsilon}}_{\text{set}}.\text{append}(\bar{\epsilon})$
end for
6. Forecast the out-of-sample set:
$Y_{\text{last}}\leftarrow Y_{\text{in-sample}}[\text{end}]$ //    Start from the last element of the in-sample set
for $i\leftarrow 1$ to $\text{length}(Y_{\text{out-of-sample}})-1$ do
     $\Delta\leftarrow\min(\Theta_{\text{set}})\cdot\min(\boldsymbol{\bar{\epsilon}}_{\text{set}})\cdot\hat{D}_{\text{out-of-sample}}[i]$ //    Compute increment
     $Y_{\text{forecast}}\leftarrow Y_{\text{last}}+\Delta$ //    Compute the adjusted naïve forecast
     $P.\text{append}(Y_{\text{forecast}})$ //    Store result
     $Y_{\text{last}}\leftarrow Y_{\text{out-of-sample}}[i]$
end for
7. Return:
return $P$
3 Simulation
This section presents a simulation study using four synthetic time series to evaluate whether the adjusted naïve forecast outperforms the original naïve forecast when movement predictions with different directional accuracy levels are provided. The performance of the proposed method under each directional accuracy level was evaluated based on the out-of-sample set of each time series.
3.1 Datasets and Setup
The dataset includes four synthetic time series, each of which represents a symmetric random walk with a certain type of variance, as summarized in Table 1. Each time series, consisting of 2,500 steps, was generated using a fixed random seed (seed = 1) in the Python environment, as shown in Figure 2. To maintain all values within the positive range, an initial positive offset of 10000 units was applied to each series. The last 500 data points were selected as the out-of-sample set, and the remaining data points were used as the in-sample set. In addition, Figure 3 shows the first-order differences of each time series to visually demonstrate the properties of the increments, and Figure 4 shows the histograms of the first-order differences for all time series, highlighting the symmetric nature of the time series.
Table 1: Variance of synthetic random walks
Variance Type Expression Parameter Range
Constant variance $\sigma_{t}^{2}=\sigma_{0}^{2}$ $\sigma_{0}^{2}=1$
Linear variance trend $\sigma_{t}^{2}=\frac{\sigma_{0}^{2}}{1+k\cdot t}$ $\sigma_{0}^{2}=1$, $k\in(0,10)$
Cyclic variance $\sigma_{t}^{2}=\sigma_{0}^{2}(1+a\cdot\sin(\omega t))$ $\sigma_{0}^{2}=1$, $\omega=\frac{2\pi}{100}$, $a\in(0,10)$
Randomly varying variance $\sigma_{t}^{2}=\sigma_{t-1}^{2}+\eta_{t},\eta_{t}\sim\mathcal{N}(0,\xi^{2})$ $\sigma_{0}^{2}=1$, $\xi^{2}\in(0,1000)$
Figure 2: Synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
Figure 3: First-order differences of four synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
Figure 4: Histograms of first-order differences of four synthetic time series: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
For each time series, the adjusted naïve forecast on the out-of-sample set was generated via Algorithm 1. The performance of each adjusted naïve forecast result was evaluated via four metrics: the RMSE, MAE, MAPE, and sMAPE. The RMSE, derived from the square root of the MSE, emphasizes larger errors, which makes it suitable for identifying significant deviations in forecasts. The MAE, on the other hand, provides an interpretable measure of the average error without overpenalizing outliers. The MAPE and sMAPE normalize the errors relative to the actual values, enabling comparisons between datasets with different scales. Together, these metrics capture both absolute and relative error characteristics, ensuring a robust and balanced evaluation. The formulas for these metrics are as follows:
$RMSE=\sqrt{\frac{1}{N}\sum\limits_{i=1}^{N}{{{({{y}_{i}}-{{{\hat{y}}}_{i}})}^{2}}}},$ (18)
$MAE=\frac{1}{N}\sum\limits_{i=1}^{N}{\left|{{y}_{i}}-{{{\hat{y}}}_{i}}\right|},$ (19)
$MAPE=\frac{100}{N}\sum\limits_{i=1}^{N}{\left|\frac{{{y}_{i}}-{{{\hat{y}}}_{i}}}{{{y}_{i}}}\right|},$ (20)
$sMAPE=\frac{100}{N}\sum\limits_{i=1}^{N}{\frac{\left|{{y}_{i}}-{{{\hat{y}}}_{i}}\right|}{\frac{\left|{{y}_{i}}\right|+\left|{{{\hat{y}}}_{i}}\right|}{2}}},$ (21)
where $N$ denotes the size of the out-of-sample set, ${{\hat{y}}_{i}}$ denotes the predicted value, and ${{y}_{i}}$ denotes the actual value.
To statistically test the performance of the adjusted naïve forecast, for each synthetic time series at a specified directional accuracy level, the process to generate the adjusted naïve forecast was repeated 100 times. Each repetition uses a unique movement prediction sequence at the same directional accuracy level, which is generated by randomly replacing a subset of actual movements in the out-of-sample set with incorrect movements. The directional accuracy level varies in two ways: initially, it changes from 0.50 to 0.56 in increments of 0.01, and then it changes from 0.55 to 1.00 in larger increments of 0.05. This approach allows for finer granularity in the lower accuracy range, where small changes in the directional accuracy can have a significant effect on the performance of the method, while using larger steps at higher accuracy levels balances the computational efficiency. Consequently, for each synthetic time series, 100 independent values were calculated for each metric at each accuracy level.
3.2 Results and Analysis
Because the directional accuracy for the out-of-sample set is predefined in each simulation process, the value of $\theta_{\text{out-of-sample}}^{*}$ can be derived directly via Eq. (14), and only the value of $\bar{\epsilon}_{\text{out-of-sample}}$ needs to be estimated. Figure 5 illustrates the distribution of the values in the $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, generated based on the sliding windows across the first-order differences of the in-sample set. The vertical reference line in Figure 5 represents the actual value of $\bar{\epsilon}_{\text{out-of-sample}}$. In the four cases, the minimum value of $\boldsymbol{\bar{\epsilon}}_{\text{set}}$ is generally an underestimate of $\bar{\epsilon}_{\text{out-of-sample}}$, except for the second case, where the synthetic time series is a random walk with a decreasing trend of linear variance. In general, the estimated value remains close to the true value. Under the assumption that the directional accuracy is consistent across the in-sample and out-of-sample sets, this estimation ensures that inequality (16) is satisfied.
Figure 5: Comparison of the true $\bar{\epsilon}_{\text{out-of-sample}}$ with $\boldsymbol{\bar{\epsilon}}_{\text{set}}$: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
Furthermore, comparisons of the two types of naïve forecasts based on the four evaluation metrics at low accuracy levels are illustrated in Figures 6, 7, 8, and 9. Each box plot in each figure represents the distribution of 100 independent values for one metric at a specific directional accuracy level, with the metric value of the original naïve forecast shown as a reference line. According to Figure 6, at low accuracy levels, such as 0.51, the RMSE of the adjusted naïve forecast is comparable to or slightly better than that of the original naïve forecast, as shown by the proximity of the median of each box plot to the reference line. As the accuracy increased beyond 0.52, most of the RMSE values fell below the baseline.
Figure 6: RMSEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 7: MAEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 8: MAPEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 9: sMAPEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Because the adjusted naïve forecast consistently outperforms the original naïve forecast at low accuracy levels in terms of the MAE, MAPE, and sMAPE, the Wilcoxon signed rank test was conducted to specifically evaluate whether the median differences between the RMSE values of the adjusted naïve forecast and those of the original naïve forecast significantly deviated from zero for accuracy levels ranging from 0.50–0.56 (Li and Johnson,, 2014). As summarized in Table 2, at a directional accuracy level of 0.51, the adjusted naïve forecast shows mixed results with respect to statistical significance. Although significant improvements were observed over the original naïve forecast for the Type 4 series (randomly varying variance, p = 0.002), the results for the Type 1 (constant variance, p = 0.915), Type 2 (linear variance trend, p = 0.781), and Type 3 (cyclic variance, p = 0.812) series were not statistically significant, indicating that the adjusted naïve forecast under an accuracy level of 0.51 may not consistently produce meaningful improvements across all types of time series. As the accuracy increased beyond 0.51, significant reductions in the RMSE (p < 0.001) were observed for all types of time series, thereby demonstrating the reliability and effectiveness of the adjusted naïve forecast at higher accuracy levels.
Table 2: RMSE results of the adjusted naïve forecast across directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Prediction Type 1 Type 2 Type 3 Type 4
RMSE Range p-value RMSE Range p-value RMSE Range p-value RMSE Range p-value
Original Naïve 0.961 - 0.00762 - 1.355 - 21.034 -
Acc = 0.50 0.961 - 0.00762 - 1.355 - 21.034 -
Acc = 0.51 0.960 $\pm$ 0.0004 0.915 0.00762 $\pm$ 0.000003 0.781 1.355 $\pm$ 0.0005 0.812 21.032 $\pm$ 0.0053 0.002
Acc = 0.52 0.960 $\pm$ 0.0008 <0.001 0.00762 $\pm$ 0.000007 <0.001 1.355 $\pm$ 0.0008 <0.001 21.024 $\pm$ 0.0102 <0.001
Acc = 0.53 0.959 $\pm$ 0.0012 <0.001 0.00761 $\pm$ 0.000010 <0.001 1.354 $\pm$ 0.0015 <0.001 21.015 $\pm$ 0.0144 <0.001
Acc = 0.54 0.959 $\pm$ 0.0017 <0.001 0.00761 $\pm$ 0.000014 <0.001 1.353 $\pm$ 0.0020 <0.001 20.995 $\pm$ 0.0199 <0.001
Acc = 0.55 0.957 $\pm$ 0.0020 <0.001 0.00760 $\pm$ 0.000020 <0.001 1.352 $\pm$ 0.0028 <0.001 20.979 $\pm$ 0.0244 <0.001
Acc = 0.56 0.956 $\pm$ 0.0023 <0.001 0.00758 $\pm$ 0.000021 <0.001 1.351 $\pm$ 0.0032 <0.001 20.957 $\pm$ 0.0278 <0.001
According to Figures 10, 11, 12, and 13, at higher accuracy levels ranging from 0.55–1.00, the adjusted naïve forecast consistently outperforms the original naïve forecast in terms of the four metrics. In particular, the improvement over the original naïve forecast becomes substantial as the directional accuracy improves. This finding demonstrates that the proposed method can effectively translate highly accurate movement predictions into improved point forecasts in the context of forecasting random walk-like time series. However, the improvement over the original naïve forecast has a limit. The proposed method allows a random walk-like series to be predicted to a certain extent, but the predicted value cannot be infinitely close to the actual value.
Figure 10: RMSEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 11: MAEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 12: MAPEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 13: sMAPEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
In general, when the directional accuracy was 0.5, the adjusted naïve forecast did not provide an improvement over the original naïve forecast. As the directional accuracy increases, the performance of the adjusted naïve forecast improves. The simulation results also demonstrate that the effectiveness of the proposed method is unaffected by the variance type of the synthetic time series, as evidenced by the consistent performance of the adjusted naïve forecast across all synthetic series. This result is consistent with the theoretical analysis of the proposed method, which highlights its robustness and potential applicability to a wide range of time series that exhibit symmetric random walk characteristics, including those with homoscedasticity as well as high-volatility scenarios.
4 Discussion
This section discusses the implications of the findings and potential limitations of this study. First, this study demonstrates an indirect regression approach that leverages movement predictions in conjunction with naïve forecasts for time series exhibiting random walk-like behavior. The theoretical analysis and simulation results suggest that accurate movement prediction can be systematically translated into improved point forecasts. This highlights the potential value of incorporating directional information into simple baseline models and suggests that enhancing movement prediction may, in some contexts, be more effective than developing increasingly complex direct regression models.
Second, in many real-world forecasting scenarios, decision-making often involves integrating predictions with heterogeneous data types, such as combining binary classifications and continuous regressions. However, such integration can be technically challenging. The proposed method provides a practical means of converting binary movement predictions into continuous point forecasts for random walk-like time series, thereby enabling more informed decision-making in applied forecasting contexts.
Third, the goal of this study is to examine the theoretical possibility of improving forecasts for random walk-like time series by incorporating directional signals rather than disputing the broader validity of the naïve forecast as the optimal benchmark. While the simulation implies that even weak directional signals can be exploited to enhance point forecasts, these improvements can be modest and context dependent. The adjusted naïve forecast serves as a second-stage method only after movement prediction is provided. Moreover, such improvements do not necessarily translate into economic gains, particularly when real-world constraints such as transaction costs, slippage, and market frictions are considered.
Finally, because the focus of this study is on the theoretical impact of directional information rather than on modeling movement predictions themselves, we rely on controlled simulations rather than empirical data. The effectiveness of the proposed adjustment depends on the consistency of the directional accuracy across different segments of the series. In practice, movement prediction models susceptible to overfitting may suffer from deteriorated performance on out-of-sample data, thereby limiting the actual benefit of incorporating such predictions into the original naïve forecast.
5 Conclusion
This study introduces a movement-prediction-adjusted naïve forecast, which is a linear combination of movement prediction and the original naïve forecast, in the challenging context of forecasting time series that exhibit symmetric random walk characteristics. The weight of the movement term is determined by two parameters derived from the in-sample data: one based on the directional accuracy of the movement prediction and the other on the mean absolute increment of the target series. The theoretical analysis of the proposed method explains how accurate movement prediction can be exploited to enhance point forecasting for a random walk-like series. It also defines a boundary criterion stating that the adjusted naïve forecast outperforms the original naïve forecast when the product of the actual optimal parameter values exceeds half the product of their estimated values. Furthermore, the simulation results demonstrate that the adjusted naïve forecast can achieve statistically significant improvements over the original naïve forecast, even with directional accuracies slightly greater than 0.50. These findings highlight the significant potential of using movement predictions to address the inherent challenges of point forecasting in a random walk-like time series.
There are several avenues for further research on the foundation of this study. First, while this research focuses on time series with symmetric random walk characteristics, future work could explore applying this movement prediction-based method to time series with nonsymmetric random walk properties, potentially broadening its applicability. Second, it is worth exploring whether the proposed forecasting method can be generalized beyond the one-dimensional case, as many real-world variables exhibit random walk behavior in a high-dimensional space. Moreover, the proposed method uses a fixed increment for simplicity. However, adapting this increment based on the availability of new data may provide more dynamic and accurate forecasts. Investigating how the increments can be adapted and whether such adaptive increments can further refine the forecasting accuracy represents a promising direction for future research.
Acknowledgments
This study did not receive any specific grants from funding agencies in the public, commercial, or not-for-profit sectors.
Appendix A. Algebraic Derivation of $\Delta MSE_{\text{in-sample}}$
According to Eq. (2), which presents the actual value of the target variable, the MSE expression for the original naïve forecast can be expressed as:
$MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (A1)
$=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t-1}+D_{t}\cdot|\epsilon_{t}|-y_{t-1}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\ |\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}\ |\epsilon_{t}|^{2}\right)$
Similarly, by substituting Eq. (2) for actual values and Eq. (4) for predicted values, the MSE of the adjusted naïve forecast can be expressed as follows:
$MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$ (A2)
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(y_{t}-\hat{y}_{t}\right)^{2}+\sum_{t\in\text{incorrect}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left[(y_{t-1}+D_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{D}_{t}\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}})\right]^{2}\right.$
$\quad\left.+\sum_{t\in\text{incorrect}}\left[(y_{t-1}+D_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{D}_{t}\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}})\right]^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right),$
Hence, the MSE difference between the adjusted naïve forecast and the original naïve forecast on the in-sample set can be derived via Eq. (A1) and Eq. (A2):
$\Delta\text{MSE}_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (A3)
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
$\quad-\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
$\quad\left.-\sum_{t\in\text{incorrect}}\left(2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
Based on the law of large numbers (LLN), when both $n_{\text{in-sample}}^{\text{correct}}$ and $n_{\text{in-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}}$ (A4)
Therefore, Eq. (A3) can be simplified as follows::
$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (A5)
$\approx\frac{1}{N}\left(2\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}})-n_{\text{in-sample}}^{\text{correct}}\cdot\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right.$
$\quad\left.-2\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}})-n_{\text{in-sample}}^{\text{incorrect}}\cdot\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)$
$=(4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}})\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$
Appendix B. Algebraic Derivation of $\Delta MSE_{\text{out-of-sample}}$
When using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as parameters for the adjusted naïve forecast on the out-of-sample set, the difference in MSE between the adjusted and original naïve forecasts on the out-of-sample set can be expressed as follows:
$\Delta\text{MSE}_{\text{out-of-sample}}$ $=\text{MSE}_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (B1)
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
$\quad-\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right.$
$\quad\left.-\sum_{t\in\text{incorrect}}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
According to the LLN, when both $m_{\text{out-of-sample}}^{\text{correct}}$ and $m_{\text{out-of-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}}$ (B2)
Therefore, Eq. (B1) can further be expressed as:
$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (B3)
$\approx\frac{1}{M}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}})-m_{\text{out-of-sample}}^{\text{correct}}\cdot\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right.$
$\quad\left.-2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}})-m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)$
$=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot ACC_{\text{out-of-sample}}\cdot\bar{\epsilon}_{\text{out-of-sample}}-2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}$
$\quad\cdot(1-ACC_{\text{out-of-sample}})\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$
$=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$