### v9

# Movement-Prediction-Adjusted Naive Forecast

Abstract
In financial time series forecasting, surpassing the naive forecast is challenging due to the randomness in the data. To address this challenge, this study proposes a novel forecast combination method, the movement-prediction-adjusted naive forecast (MPANF), which is designed to improve point forecasts beyond the naive baseline. Specifically, MPANF integrates two forecasting components: a naive forecast and a movement prediction. The final forecast is generated by adjusting the naive forecast with a movement prediction term, the weight of which is the product of two in-sample quantities: one is a coefficient determined from the movement prediction accuracy and the other is the mean absolute increment. The performance of MPANF was evaluated on eight financial time series via standard metrics, including the RMSE, MAE, MAPE, and sMAPE. Under modest movement prediction accuracy slightly above 0.55, MPANF generally outperforms common benchmarks such as the naive forecast, naive forecast with drift, integrated moving average of order (1,1) (IMA(1,1)), and linear regression. These findings suggest that MPANF can serve as an effective second-stage method when reliable movement predictions are available.
Keywords Financial time series  $\cdot$ Point forecasting  $\cdot$ Movement prediction  $\cdot$ Naive forecast  $\cdot$ Second-stage forecast combination
1 Introduction
Financial time series forecasting plays a crucial role in many decision-making processes. The widespread demand for accurate point forecasts in the financial market has attracted considerable attention from researchers and practitioners. However, although financial time series have been extensively studied as test cases for forecasting theories and methods, randomness in the data often poses significant challenges in generating accurate forecasts (Hamilton,, 2020). One prominent theory on this challenge is the efficient market hypothesis (EMH), which suggests that asset prices incorporate all available information, rendering historical prices unpredictable (Fama,, 1970). The naive forecast, or persistent forecast, which assumes that the next value is equal to the current value, is frequently regarded as the optimal benchmark for these tasks.
In support of the EMH, Moosa, (2013) reported that as the volatility of financial time series increases, the root mean square error (RMSE) of the forecasting models increases more quickly than that of the naive forecast. Moreover, Moosa and Burns, (2014) argued that claims of outperforming the naive forecast in terms of the RMSE are often misleading, frequently involving introduced dynamics without rigorous statistical validation. Despite numerous attempts to exploit weak structural signals in financial time series, existing forecasting methods—including statistical, machine learning, and deep learning approaches—have typically failed to outperform naive forecasts (Kilian and Taylor,, 2003; Moosa and Burns,, 2016; Thakkar and Chaudhari,, 2021; Petropoulos et al.,, 2022; Ellwanger and Snudden,, 2023; Hewamalage et al.,, 2023; Zeng et al.,, 2023; Beck et al.,, 2025).
From the forecaster’s perspective, the unpredictability of a variable could be attributed to incomplete knowledge of the underlying processes controlling the data, ignorance of the various contributing factors, and imperfections in analytical methods (Laplace,, 2012). In other words, randomness in data is not inherent but rather apparent (Crutchfield and Feldman,, 2003). Although useful information may be present in the input space, it can be overwhelmed by irrelevant information, making it difficult for the regression method to exploit the useful content effectively (John et al.,, 1994; Guyon and Elisseeff,, 2003; Verleysen and François,, 2005). Moreover, identifying the predictors on which the target variable truly depends is a complex task (Granger,, 1969; Stock and Watson,, 2002). Nevertheless, these theories and empirical evidence reveal only the practical difficulty of finding exploitable patterns; they do not eliminate the possibility of discovering any predictability in the data.
In a forecasting task, if the mutual information between the predictors and the target variable is greater than zero, it indicates that a dependency between them exists (Kraskov et al.,, 2004; Cover and Thomas,, 2012). In practical terms, this means that the predictors can reduce uncertainty in binary forecasting tasks, such as determining whether the target value will rise or fall in future time steps. Although this does not guarantee high predictive accuracy, it suggests that, theoretically, one can build a model that performs better than random guessing (Ircio et al.,, 2020). In contrast, for a point forecasting task, the continuous nature of the target means that even if a small gain in mutual information exists, it only marginally reduces the overall uncertainty and does not sufficiently narrow down the precise numerical output (Beraha et al.,, 2019). Given the difficulty of generating precise numerical predictions from noisy data in financial time series, researchers have increasingly shifted their focus from point forecasting to movement prediction.
Indeed, movement prediction is often considered easier than point forecasting in financial time series (Taylor,, 2008; Moosa,, 2013). Behavioral finance studies further suggest that investor sentiment and psychological factors can generate predictable movement patterns (De Bondt and Thaler,, 1985; Lee et al.,, 1991; Baker and Wurgler,, 2007; Barberis et al.,, 1998). Consequently, predictors such as sentiment indicators are often used to improve the directional accuracy in these movement prediction tasks. For example, Weng et al., (2017) incorporated Google news counts and Wikipedia page views alongside historical data to predict the next day’s movement of AAPL stock via support vector machines (SVMs), achieving up to 0.80 directional accuracy. Ma et al., (2023) proposed a multisource aggregated classification (MAC) method for the prediction of stock price movement, incorporating the numerical features and market-driven news sentiments of target stocks, and the directional accuracy reached 0.70. Wang et al., (2025) used multimodal data that included news data to train a hybrid model consisting of a long short memory network (LSTM) and a transformer for the prediction of stock movement, and the directional accuracy fell within the range of 0.57–0.87. Depending on the complexity of the markets and availability of data, the movement prediction accuracy of financial time series typically ranges between 0.55 and 0.80 (Bustos and Pomares-Quimbaya,, 2020).
These accuracy levels, which are consistently above the 0.50 threshold of random guessing, confirm that financial time series movement prediction is feasible. This naturally raises the question of whether the directional information captured by movement prediction can be systematically incorporated into point forecasting to increase accuracy. A forecasting framework for this purpose should align with the “less is more” principle, which posits that reducing informational complexity can enhance decision-making (Gigerenzer and Brighton,, 2009). Accordingly, it should rely solely on the naive forecast and the movement prediction as inputs, thereby avoiding additional noisy predictors (see Figure 1). The naive forecast, which represents the current value, is most closely related to the future value, whereas the movement prediction contributes new directional information not contained within the naive forecast. Unfortunately, previous studies either focused exclusively on movement prediction or applied regression directly to noisy data. Although forecast combination and consensus forecasting are well established, they typically blend multiple point forecasts rather than use directional signals to improve point forecast accuracy. The use of movement prediction alone for this purpose remains unexplored, leaving a notable gap in systematic approaches for integrating movement prediction outcomes with naive forecasts in time series forecasting.
Figure 1: Direct and indirect frameworks for point forecasting.
To investigate the use of movement prediction for improving point forecasting in financial time series, we propose the movement-prediction-adjusted naive forecast (MPANF) — a forecast combination method in which the next-step point forecast is obtained by adjusting the original naive forecast with a movement prediction term. The weight assigned to this term is computed as the product of two in-sample quantities: a coefficient determined by the movement prediction accuracy and the mean absolute increment.
The effectiveness of the proposed method was evaluated through experiments on eight real-world financial time series from the U.S. market. For all the experiments, the naive, naive with drift, integrated moving average of order (1,1) (IMA (1,1)), and linear regression serve as the baselines, with the RMSE, mean absolute error (MAE), mean absolute percentage error (MAPE) and symmetric mean absolute percentage error (sMAPE) used as evaluation metrics.
This study makes three significant contributions to the literature:
  • First, it proposes a novel forecast combination method that explicitly integrates movement predictions into point forecasting, uniquely linking these two distinct forecasting tasks. Although binary classification to guide regression is not new, exclusively using the target variable’s own movement prediction to inform its point forecast offers a fresh methodological perspective in forecasting.
  • Second, it offers the first theoretical justification and mathematical analysis demonstrating how and why movement predictions alone can be exploited to generate point forecasts that outperform the naive forecast. This theoretical foundation highlights the informational value embedded in directional signals and distinguishes the proposed method from prior heuristic or empirical approaches.
  • Third, it provides empirical insights into the conditional predictability of financial time series, demonstrating that series traditionally considered unpredictable can exhibit meaningful predictability when accurate movement prediction is available.
The remainder of this paper is organized as follows. Section 2 introduces the details of the MPANF and explains how to derive it and under what conditions it can outperform the naive forecast. Section 3 presents empirical results on the performance of the MPANF via real-world financial time series. Section 4 discusses the implications of the findings and limitations of this study. Finally, section 5 concludes the paper and outlines future research directions.
2 Method
This section presents the details of the MPANF, which aims to outperform the naive forecast in point forecasting by incorporating the target variable’s movement prediction. We begin by deriving the general formulation of the MPANF. Subsequently, we optimize the necessary coefficient using the in-sample set and conduct uncertainty analysis for out-of-sample forecasts.
2.1 General Formulation of the MPANF
The value of a time series at time step $t$ can be expressed as the sum of its previous value at time step t1t-1 and an increment ${\epsilon}_{t}$}, which represents the value change between two adjacent observations. This relationship can be expressed as:
${{y}_{t}}={{y}_{t-1}}+{{\epsilon}_{t}},$ (1)
where ${y}_{t}$ is the point value at time step $t$ and ${y}_{t-1}$ is the point value at time step t1t-1. The increment $\epsilon_{t}$ at time step t can be expressed as a product of two terms: a nonnegative term $|\epsilon_{t}|$ representing the magnitude of the increment and a sign term ${D}_{t}$, which indicates the direction of the increment between two adjacent points and takes a value of +1+1 for upward movement and 1-1 for downward movement. This allows ${y}_{t}$ to be reformulated as follows:
${{y}_{t}}={{y}_{t-1}}+{{D}_{t}}\cdot|\epsilon_{t}|;|\epsilon_{t}|\geq 0$ (2)
The nonnegative term $|\epsilon_{t}|$ can be further expressed as a product of two terms:
${{y}_{t}}={{y}_{t-1}}+{{D}_{t}}\cdot{\theta_{t}}\cdot\bar{\epsilon};{\theta_{t}}\geq 0,\bar{\epsilon}>0,$ (3)
where $\bar{\epsilon}$ represents the mean absolute increment of the time series, which can be interpreted as a base magnitude of the increments. The term $\theta_{t}$ is introduced as a nonnegative coefficient that controls the magnitude of the increment at time step $t$. Eq. (3) implies that the point forecast of the target variable can be generated if ${D}_{t}$} can be predicted and $\theta_{t}$ can be estimated. Consider a simple forecasting strategy in which the increments for each future time step are set to have a fixed magnitude; in other words, we set $\theta_{t}$ to be a constant value and derive the one-step-ahead point forecast at time step $t$, denoted as $\hat{y}_{t}$, as follows:
${{\hat{y}}_{t}}={{y}_{t-1}}+\hat{D}_{t}\cdot\theta\cdot\bar{\epsilon};\bar{\epsilon}>0,$ (4)
where ${\hat{D}}_{t}$} is the movement prediction, $\theta$ is a fixed scalar, and $\theta\in\mathbb{R}$. Therefore, the point forecast given by Eq. (4) is an adjusted naive forecast, where the adjustment is a movement prediction term weighted by the product of $\theta$ and $\bar{\epsilon}$. Hence, it is termed the movement-prediction-adjusted naive forecast (MPANF).
2.2 Determination of the Optimal Coefficient $\theta$
Since $\bar{\epsilon}$ in Eq. (4) is computed as the mean absolute increment of the in-sample set, only the coefficient $\theta$ needs to be optimized. Suppose that $\theta$ is optimized by minimizing the in-sample mean squared error (MSE) of the MPANF. Since the in-sample MSE of the original naive forecast is constant, this optimization is equivalent to maximizing the in-sample MSE difference between the original naive forecast and the MPANF:
$\min MSE_{\text{in-sample}}^{\text{MPANF}}\equiv\max\left(MSE_{\text{in-sample}}^{\text{Naive}}-MSE_{\text{in-sample}}^{\text{MPANF}}\right),$ (5)
where $MSE_{\text{in-sample}}^{\text{Naive}}$ and $MSE_{\text{in-sample}}^{\text{MPANF}}$ denote the in-sample MSE of the original naive forecast and that of the MPANF, respectively. Suppose the in-sample set has $N$ time steps, and the movement prediction yields $n_{\text{in-sample}}^{\text{correct}}$ correct and $n_{\text{in-sample}}^{\text{incorrect}}$ incorrect directions; then, the in-sample movement prediction accuracy, denoted as $ACC_{\text{in-sample}}$, can be expressed as follows:
$ACC_{\text{in-sample}}=\frac{n_{\text{in-sample}}^{\text{correct}}}{n_{\text{in-sample}}^{\text{incorrect}}+n_{\text{in-sample}}^{\text{correct}}}$ (6)
Accordingly, the MSE difference between the naive forecast and the MPANF, denoted as $\Delta MSE_{\text{in-sample}}$, is expressed as follows:
$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{Naive}}-MSE_{\text{in-sample}}^{\text{MPANF}}$ (7)
$=(4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}})\cdot\bar{\epsilon}_{\text{in-sample}}^{2},$
where $\theta_{\text{in-sample}}$ denotes the coefficient $\theta$ used by the MPANF on the in-sample set and $\bar{\epsilon}_{\text{in-sample}}$ denotes the mean of $|\epsilon_{t}|$ calculated over the in-sample set. A detailed algebraic derivation of the $\Delta MSE_{\text{in-sample}}$ is provided in Appendix A.
According to Eq. (7), maximizing $\Delta MSE_{\text{in-sample}}$ can further be reduced to maximizing the following function of $\theta_{\text{in-sample}}$:
$f(\theta_{\text{in-sample}})=4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}},$ (8)
Since Eq. (8) is a concave function, differentiating it with respect to $\theta_{\text{in-sample}}$ and setting the derivative to zero yields the optimal solution:
Proposition 1 (Optimal coefficient). The optimal in-sample coefficient $\theta_{\text{in-sample}}^{*}$, which maximizes the in-sample MSE difference between the naive forecast and the MPANF, is given by:
$\theta_{\text{in-sample}}^{*}=2\cdot ACC_{\text{in-sample}}-1$ (9)
This equation indicates that the optimal coefficient $\theta$ depends solely on the in-sample movement prediction accuracy and is independent of the specific movement predictions or the characteristics of the time series.
Substituting Eq. (9) back into Eq. (7) yields:
Proposition 2 (Performance improvement). Under the optimal coefficient setting using $\theta_{\text{in-sample}}^{*}$, the MPANF achieves a nonnegative improvement over the naive forecast in terms of the in-sample MSE:
$\max\Delta MSE_{\text{in-sample}}=\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\geq 0$ (10)
Proposition 2 confirms that as the directional accuracy increases beyond 0.5, the value of $\theta_{\text{in-sample}}^{*}$ increases from 0.0, leading to progressively greater improvements over the naive forecast. At a directional accuracy of 1.0, the maximum achievable improvement is $\bar{\epsilon}_{\text{in-sample}}^{2}$. Thus, with the optimal coefficient setting, the performance of the MPANF matches or surpasses that of the naive forecast on the in-sample set:
$MSE_{\text{in-sample}}^{\text{MPANF}}\leq MSE_{\text{in-sample}}^{\text{Naive}}$ (11)
2.3 Uncertainty Analysis of Out-of-Sample Forecasts
We now examine the out-of-sample performance of the MPANF. Suppose that the movement predictions on the out-of-sample set with $M$ time steps yield $m_{\text{out-of-sample}}^{\text{correct}}$ correct and $m_{\text{out-of-sample}}^{\text{incorrect}}$ incorrect predictions; then, the out-of-sample movement prediction accuracy, denoted as $ACC_{\text{out-of-sample}}$, can be expressed as:
$ACC_{\text{out-of-sample}}=\frac{m_{\text{out-of-sample}}^{\text{correct}}}{m_{\text{out-of-sample}}^{\text{incorrect}}+m_{\text{out-of-sample}}^{\text{correct}}}$ (12)
Suppose that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are used by the MPANF on the out-of-sample set; then, the out-of-sample MSE difference between the naive forecast and the MPANF, denoted as $\Delta MSE_{\text{out-of-sample}}$, is expressed as:
$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{Naive}}-MSE_{\text{out-of-sample}}^{\text{MPANF}}$ (13)
$=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2},$
where $\bar{\epsilon}_{\text{out-of-sample}}$ denotes the out-of-sample mean absolute increment and $\theta_{\text{out-of-sample}}^{*}$ denotes the optimal out-of-sample coefficient $\theta$, which is derived according to Proposition 1:
$\theta_{\text{out-of-sample}}^{*}=2\cdot ACC_{\text{out-of-sample}}-1.$ (14)
A detailed algebraic expansion of the $\Delta MSE_{\text{out-of-sample}}$ is provided in Appendix B. For the MPANF to outperform the naive forecast, the $\Delta MSE_{\text{out-of-sample}}$ should be no less than zero. Therefore, according to Eq. (13), the following inequality must hold:
$\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}\geq\frac{1}{2}\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}$ (15)
Given that $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are the estimates of $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$, respectively, inequality (15) can be generalized as:
Proposition 3 (Out-of-sample performance criterion). The necessary condition for the MPANF to achieve an improvement over the naive forecast on the out-of-sample set in terms of the MSE is as follows:
$\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}\geq\frac{1}{2}\cdot\hat{\theta}_{\text{out-of-sample}}\cdot\hat{\bar{\epsilon}}_{\text{out-of-sample}},$ (16)
where $\hat{\theta}_{\text{out-of-sample}}$ denotes the estimate of $\theta_{\text{out-of-sample}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ denotes the estimate of $\bar{\epsilon}_{\text{out-of-sample}}$.
Proposition 3 establishes a boundary criterion when the MPANF can outperform the naive forecast. We can treat $\theta_{\text{out-of-sample}}$ and $\bar{\epsilon}_{\text{out-of-sample}}$ as two random variables, each following a distribution defined over nonnegative values. Accordingly, inequality (16) represents a probabilistic event whose likelihood depends on the realized values of these two variables. Smaller values of these two variables generally increase the likelihood of improvement. However, as shown in Eq. (13), the resulting gains may be marginal. The MPANF can realize its maximum potential improvement only when $\hat{\theta}_{\text{out-of-sample}}$ and $\hat{\bar{\epsilon}}_{\text{out-of-sample}}$ are close to their respective true values, $\theta_{\text{out-of-sample}}^{*}$ and $\bar{\epsilon}_{\text{out-of-sample}}$.
In practice, the MPANF on the out-of-sample set can be derived via Algorithm 1.
Algorithm 1 Movement-prediction–adjusted naive forecast on the out-of-sample set
1:Inputs:
[…]
6:Output:
7:$P$ $\triangleright$ Final point forecasts
8:Initialize:
[…]
10:Step 1: Determine $\theta$ and $\bar{\epsilon}$ from in-sample data
[…]
16:Step 2: Generate forecasts on the out-of-sample set
[…]
18:for $i\leftarrow 1,\dots,\text{length}(Y_{\text{out-of-sample}})$ do
19:  if $\hat{D}_{\text{out-of-sample}}[i]=1$ then
20:   $\Delta\leftarrow\theta\cdot\bar{\epsilon}$ $\triangleright$ Upward
21:  else
22:   $\Delta\leftarrow-\theta\cdot\bar{\epsilon}$ $\triangleright$ Downward
23:  end if
24:  $\hat{Y}_{i}\leftarrow Y_{\text{last}}+\Delta$
25:  $P.\text{append}(\hat{Y}_{i})$
26:  $Y_{\text{last}}\leftarrow Y_{\text{out-of-sample}}[i]$
27:end for
28:return $P$
3 Empirical Results
This section presents the empirical evaluation of the MPANF using real-world data. It includes a description of the datasets, details of the experimental setup, and a performance comparison of the proposed method against several baselines.
3.1 Datasets and Setup
In the empirical analysis, we first generate movement predictions for each target series and then apply the MPANF to produce the final point forecasts. Since developing a complex movement prediction method is beyond the scope of this study, the movement predictions are obtained using a simple co-movement–based approach with a selected exogenous variable, as detailed in Appendix C. This two-step design reflects the nature of the MPANF as a second-stage method that relies on accurate and consistent directional inputs.
Given the availability of data and the well-documented strong co-movement between the Financial Times Stock Exchange (FTSE) index and U.S. stock prices (Sarwar,, 2020; Gupta et al.,, 2024), together with the fact that the FTSE open price is recorded approximately 13 hours before the U.S. market close price on the same date, we selected eight U.S. financial time series as the target series, including Apple (AAPL), Boeing (BA), Google (GOOGL), Goldman Sachs (GS), Halliburton (HAL), Brent crude oil (OIL), Schlumberger (SLB), and Tesla (TSLA), based on their strong predictive alignment with FTSE movements. Accordingly, the FTSE open price is used as the exogenous variable for generating movement predictions.
Historical data for the FTSE open price and the close prices of the eight target series were obtained from Yahoo Finance, covering the period from October 8, 2014, to September 13, 2024. The FTSE data were aligned with each target series on a day-by-day basis, with extra observations removed and missing values imputed via forward filling (Junninen et al.,, 2004). After alignment, all series were standardized to a uniform length of 2,500 observations and then split into in-sample and out-of-sample sets at a 1:1 ratio, ensuring sufficient data for both model estimation and evaluation.The close prices of the eight target variables and the FTSE open price are illustrated in Figures 2 and 3, respectively, with a brief description provided in Table 1. The in-sample movement prediction accuracies for the eight target series all exceeded 0.55, indicating that the directional signals derived from the FTSE open price were meaningful.
Figure 2: Time series of eight target variables.
Figure 3: Open price of FTSE.
Table 1: Statistics of nine financial time series.
Time series Count Mean Std. Dev. Minimum Median Maximum $\mathbf{ACC}_{\text{in-sample}}$ $\mathbf{\bar{\epsilon}}_{\text{in-sample}}$
AAPL (Close price) 2500 88.46 62.00 20.72 54.16 234.55 0.5536 0.3842
BA (Close price) 2500 211.21 80.90 95.01 195.92 430.30 0.5568 2.5867
GOOGL (Close price) 2500 78.45 41.52 24.79 61.44 190.93 0.5504 0.4624
GS (Close price) 2500 239.35 89.70 117.58 202.27 510.25 0.5504 1.9030
HAL (Close price) 2500 30.99 9.61 4.32 33.30 151.09 0.5680 0.5252
OIL (Close price) 2500 66.43 18.59 19.33 65.20 127.98 0.5576 0.9067
SLB (Close price) 2500 45.94 15.45 11.02 49.49 74.91 0.5536 0.6596
TSLA (Close price) 2500 107.38 109.43 9.58 23.72 409.97 0.5536 0.3534
FTSE (Open price) 2500 7088.28 589.13 4993.90 7210.00 8445.80
Since many sophisticated models, such as machine learning or deep learning-based models, struggle to provide significant improvements over naive forecasts in financial time series forecasting, simple but strong baselines, such as naive, naive with drift, IMA (1,1), and linear regression, are selected for performance comparison Rogers, (1995); Jadevicius and Huston, (2015); Franses, (2020). In particular, the linear regression model, which acts as a meta-learner to combine the naive forecast and the movement prediction, uses only the current observed value and the predicted movement as predictors.
The performance of each method was evaluated via four metrics: the RMSE, MAE, MAPE, and sMAPE. The RMSE, derived from the square root of the MSE, emphasizes larger errors, which makes it suitable for identifying significant deviations in forecasts. The MAE, on the other hand, provides an interpretable measure of the average error without over penalizing outliers. The MAPE and sMAPE normalize the errors relative to the actual values, enabling comparisons between datasets with different scales. Together, these metrics capture both absolute and relative error characteristics, ensuring a robust and balanced evaluation. The formulas for these metrics are as follows:
$RMSE=\sqrt{\frac{1}{N}\sum\limits_{i=1}^{N}{{{({{y}_{i}}-{{{\hat{y}}}_{i}})}^{2}}}},$ (17)
$MAE=\frac{1}{N}\sum\limits_{i=1}^{N}{\left|{{y}_{i}}-{{{\hat{y}}}_{i}}\right|},$ (18)
$MAPE=\frac{100}{N}\sum\limits_{i=1}^{N}{\left|\frac{{{y}_{i}}-{{{\hat{y}}}_{i}}}{{{y}_{i}}}\right|},$ (19)
$sMAPE=\frac{100}{N}\sum\limits_{i=1}^{N}{\frac{\left|{{y}_{i}}-{{{\hat{y}}}_{i}}\right|}{\frac{\left|{{y}_{i}}\right|+\left|{{{\hat{y}}}_{i}}\right|}{2}}},$ (20)
where $N$ denotes the size of the out-of-sample set, ${\hat{y}}_{i}$} denotes the predicted value, and ${y}_{i}$} denotes the actual value.
3.2 Results and Analysis
The out-of-sample forecasting results for the eight financial time series are illustrated in Figure 4. Given that each out-of-sample set consists of 1250 time steps and that predictions from different methods visually overlap to a large extent, only the first five predicted points from each set are displayed for clarity. Small differences between the MPANF and the other forecasting methods are observed because the overall in-sample movement prediction accuracy for each time series is just slightly above 0.55, leading to a small coefficient $\theta$ (approximately 0.1) and a modest adjustment of the naive forecast at each time step. The naive forecast, naive forecast with drift, and IMA(1,1) forecasts yield similar predictions, reinforcing their suitability as baseline methods for time series forecasting. Moreover, because the linear regression model closely mirrors the structure of the MPANF, its performance is similarly dependent on movement prediction accuracy and therefore remains close to that of the MPANF.
Figure 4: Out-of-sample forecasts by different methods.
Tables 2, 3, 4, and 5 present performance comparisons between the MPANF and the baseline methods for eight out-of-sample sets, reported in terms of the RMSE, MAE, MAPE, and sMAPE, respectively. Notably, the MPANF consistently outperforms the naive forecast, with the only exceptions being the MAPE and sMAPE for the GOOGL dataset. Specifically, the MPANF achieves the lowest RMSE in 4 out of 8 cases, the lowest MAE in 6 out of 8 cases, the lowest MAPE in 6 out of 8 cases, and the lowest sMAPE in 5 out of 8 cases, demonstrating its overall superiority over the baseline methods.
Table 2: RMSE comparison across forecasting methods for selected stocks
Stock Naive Naive + Drift IMA(1,1) LR MPANF
AAPL 2.5937 2.5925 2.5979 2.5942 2.5931
BA 5.9080 5.9187 5.9166 5.8932 5.8829
GOOGL 2.2093 2.2085 2.2108 2.2114 2.2088
GS 5.3671 5.3660 5.3727 5.4039 5.3485
HAL 0.7334 0.7341 0.7332 0.7261 0.7271
OIL 1.8909 1.8912 1.8976 1.8940 1.8864
SLB 0.9401 0.9410 0.9401 0.9331 0.9343
TSLA 8.1524 8.1525 8.1677 8.3351 8.1515
Table 3: MAE comparison across forecasting methods for selected stocks
Stock Naive Naive + Drift IMA(1,1) LR MPANF
AAPL 1.8950 1.8935 1.8975 1.8958 1.8943
BA 3.9691 3.9821 3.9750 3.9657 3.9533
GOOGL 1.5680 1.5660 1.5727 1.5746 1.5674
GS 3.9743 3.9737 3.9811 4.0021 3.9534
HAL 0.5423 0.5429 0.5439 0.5376 0.5368
OIL 1.2959 1.2981 1.3000 1.3073 1.2903
SLB 0.6901 0.6910 0.6907 0.6852 0.6843
TSLA 5.5732 5.5733 5.5775 5.8155 5.5722
Table 4: MAPE (%) comparison across forecasting methods for selected stocks
Stock Naive Naive + Drift IMA(1,1) LR MPANF
AAPL 1.4081 1.4069 1.4065 1.4079 1.4074
BA 2.0601 2.0675 2.0633 2.0585 2.0514
GOOGL 1.4357 1.4340 1.4380 1.4406 1.4359
GS 1.4037 1.4036 1.4046 1.4074 1.3967
HAL 2.3517 2.3532 2.3667 2.3277 2.3216
OIL 1.8379 1.8409 1.8430 1.8456 1.8296
SLB 2.1522 2.1532 2.1574 2.1355 2.1291
TSLA 2.9511 2.9513 2.9549 3.0534 2.9509
Table 5: sMAPE (%) comparison across forecasting methods for selected stocks
Stock Naive Naive + Drift IMA(1,1) LR MPANF
AAPL 1.4088 1.4072 1.4072 1.4088 1.4081
BA 2.0542 2.0591 2.0591 2.0500 2.0457
GOOGL 1.4352 1.4331 1.4375 1.4410 1.4354
GS 1.4035 1.4032 1.4045 1.4090 1.3966
HAL 2.3377 2.3427 2.3537 2.3051 2.3084
OIL 1.8290 1.8328 1.8336 1.8377 1.8210
SLB 2.1465 2.1503 2.1521 2.1260 2.1240
TSLA 2.9518 2.9520 2.9554 3.0767 2.9515
In addition, baseline methods such as naive with drift, IMA(1,1), or linear regression occasionally yield lower RMSE values (as observed for AAPL, GOOGL, HAL, and OIL), lower MAE values (AAPL, GOOGL), lower MAPE values (AAPL, GOOGL), and lower sMAPE values (AAPL, GOOGL, and HAL). However, the performances of these methods are inconsistent; they frequently fall short of the naive forecast in other cases.
Notably, although linear regression also uses movement prediction as an input feature and may occasionally outperform the naive forecast, it lacks a theoretical explanation for such improvement. Moreover, its performance generally lags behind that of the MPANF. This performance gap arises because linear regression is more susceptible to overfitting. While it closely fits the in-sample data, its out-of-sample performance can deteriorate significantly if the movement prediction accuracy decreases. In contrast, the coefficient determination in the MPANF is guided by theoretically derived optimal conditions rather than by data-driven minimization of in-sample errors. This design inherently introduces regularization, effectively mitigating overfitting and enhancing the method’s generalization ability.
Additionally, Figures 5, 6, 7, and 8 present the performance of each forecasting method across the four evaluation metrics. Overall, the MPANF generates more accurate and robust forecasts than the baseline methods do and consistently improves upon the naive forecast across diverse financial forecasting scenarios. These empirical results are consistent with the theoretical analysis, underscoring the method’s broad applicability to financial time series forecasting, including series characterized by heteroscedasticity and high volatility.
Figure 5: Comparison of the RMSEs of different methods.
Figure 6: Comparison of the MAEs of different methods.
Figure 7: Comparison of the MAPEs of different methods.
Figure 8: Comparison of the sMAPEs of different methods.
4 Discussion
This section discusses the implications of the findings and the potential limitations of the study.
First, the results demonstrate the effectiveness of a second-stage forecasting method that combines movement predictions with naive forecasts for financial time series forecasting. Both the theoretical analysis and the empirical evidence indicate that accurate movement prediction can be systematically translated into improved point forecasts. This underscores the value of incorporating directional information into a naive forecast and suggests that, in some contexts, the indirect forecasting approach may be more effective than developing complex regression models.
Second, in many real-world forecasting applications, decision-making requires integrating predictions from heterogeneous data types, such as binary classifications and continuous regressions. Such integration can be technically challenging. The proposed method offers a practical solution by converting binary movement predictions into continuous point forecasts, thereby facilitating more informed decisions in forecasting scenarios.
Third, while the empirical results show that even relatively weak directional signals can enhance point forecasts, the magnitude of such improvements is often modest and context dependent. Moreover, these gains do not necessarily translate into economic benefits, especially when real-world factors such as transaction costs, slippage, and market frictions are taken into account.
Finally, although the MPANF imposes no restrictions on the choice of movement prediction models, its effectiveness depends critically on the consistency and reliability of directional accuracy across different segments of the dataset. In practice, predicting the movements of financial time series remains challenging, and models prone to overfitting may suffer substantial degradation in out-of-sample performance. Such deterioration can reduce, or even eliminate, the practical advantages of incorporating movement predictions into the naive forecast. In other words, the applicability of the MPANF is conditional on the availability of a meaningful and robust movement prediction for the target variable.
5 Conclusions
This study introduces the MPANF, a forecast combination method that adjusts the naive forecast using the movement prediction of the target variable. The approach is particularly effective for time series in which surpassing the naive forecast is typically difficult but accurate movement prediction is available. The weight assigned to the movement term is calculated as the product of a coefficient derived from the in-sample movement prediction accuracy and the in-sample mean absolute increment. Theoretical analysis explains why this combination can yield point forecasts that are superior to the naive forecast. Empirical evaluation on eight financial time series confirms the MPANF’s ability to achieve robust improvements over the naive forecast, even when the directional accuracy is only modestly above 0.55. Together, these findings underscore the potential of leveraging movement predictions to enhance point forecasts in challenging forecasting scenarios.
Several avenues for future research emerge from this work. First, because the present evaluation was conducted under modest directional accuracy, developing more advanced movement prediction models and assessing MPANF performance at higher accuracy levels could enhance its practical value in real-world trading. Second, applying MPANF to other time series domains—particularly those where the naive forecast is hard to beat yet movement prediction is feasible—could broaden its applicability. Third, extending the MPANF beyond the univariate case would address the many forecasting tasks that arise in high-dimensional settings. Finally, while this study employs a fixed adjustment for simplicity, exploring adaptive adjustment schemes may further improve the forecasting accuracy.
Acknowledgments
This study did not receive any specific grants from funding agencies in the public, commercial, or not-for-profit sectors.
Appendix A. Algebraic Derivation of $\Delta MSE_{\textnormal{in-sample}}$
According to Eq. (2), the in-sample MSE expression for the naive forecast, denoted as $MSE_{\text{in-sample}}^{\text{Naive}}$, can be expressed as:
$MSE_{\text{in-sample}}^{\text{Naive}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (A1)
$=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t-1}+D_{t}\cdot|\epsilon_{t}|-y_{t-1}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\ |\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}\ |\epsilon_{t}|^{2}\right)$
By substituting Eq. (2) into $y_{t}$ and Eq. (4) into $y_{t-1}$, the in-sample MSE of the MPANF, denoted as $MSE_{\text{in-sample}}^{\text{MPANF}}$, can be expressed as follows:
$MSE_{\text{in-sample}}^{\text{MPANF}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$ (A2)
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(y_{t}-\hat{y}_{t}\right)^{2}+\sum_{t\in\text{incorrect}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left[(y_{t-1}+D_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{D}_{t}\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}})\right]^{2}\right$.
$\quad\left.+\sum_{t\in\text{incorrect}}\left[(y_{t-1}+D_{t}\cdot|\epsilon_{t}|)-(y_{t-1}+\hat{D}_{t}\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}})\right]^{2}\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right$.
$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right),$
Hence, the in-sample MSE difference between the naive forecast and the MPANF, denoted as $\Delta MSE_{\text{in-sample}}$, can be derived via Eq. (A1) and Eq. (A2):
$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{Naive}}-MSE_{\text{in-sample}}^{\text{MPANF}}$ (A3)
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
$\quad-\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right$.
$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right$.
$\quad\left.-\sum_{t\in\text{incorrect}}\left(2\cdot\theta_{\text{in-sample}}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
Based on the law of large numbers (LLN), under idealized conditions where both $n_{\text{in-sample}}^{\text{correct}}$ and $n_{\text{in-sample}}^{\text{incorrect}}$ are sufficiently large, and the two sets of increments corresponding to correct and incorrect movement predictions are unbiased and representative of the overall increment set, the following approximations hold:
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}}$ (A4)
Therefore, Eq. (A3) can be simplified as follows::
$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{Naive}}-MSE_{\text{in-sample}}^{\text{MPANF}}$ (A5)
$\approx\frac{1}{N}\left(2\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(n_{\text{in-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{in-sample}})-n_{\text{in-sample}}^{\text{correct}}\cdot\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right$.
$\quad\left.-2\cdot\theta_{\text{in-sample}}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(n_{\text{in-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{in-sample}})-n_{\text{in-sample}}^{\text{incorrect}}\cdot\theta_{\text{in-sample}}^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)$
$=(4\cdot\theta_{\text{in-sample}}\cdot ACC_{\text{in-sample}}-\theta_{\text{in-sample}}^{2}-2\cdot\theta_{\text{in-sample}})\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$
Appendix B. Algebraic Derivation of $\Delta MSE_{\textnormal{out-of-sample}}$
When $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are used by the MPANF on the out-of-sample set, the out-of-sample MSE difference between the naive forecast and the MPANF, denoted as $\Delta MSE_{\text{out-of-sample}}$, can be expressed as follows:
$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{Naive}}-MSE_{\text{out-of-sample}}^{\text{MPANF}}$ (B1)
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
$\quad-\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right$.
$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right$.
$\quad\left.-\sum_{t\in\text{incorrect}}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in-sample}}+\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)\right)$
According to the LLN, under idealized conditions where both $m_{\text{in-sample}}^{\text{correct}}$ and $m_{\text{in-sample}}^{\text{incorrect}}$ are sufficiently large, and the two sets of increments corresponding to correct and incorrect movement predictions are unbiased and representative of the overall increment set, the following approximations hold:
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}}$ (B2)
Therefore, Eq. (B1) can further be expressed as:
$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{Naive}}-MSE_{\text{out-of-sample}}^{\text{MPANF}}$ (B3)
$\approx\frac{1}{M}\left(2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(m_{\text{out-of-sample}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out-of-sample}})-m_{\text{out-of-sample}}^{\text{correct}}\cdot\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right$.
$\quad\left.-2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot(m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out-of-sample}})-m_{\text{out-of-sample}}^{\text{incorrect}}\cdot\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}\right)$
$=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot ACC_{\text{out-of-sample}}\cdot\bar{\epsilon}_{\text{out-of-sample}}-2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}$
$\quad\cdot(1-ACC_{\text{out-of-sample}})\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$
$=2\cdot\theta_{\text{in-sample}}^{*}\cdot\bar{\epsilon}_{\text{in-sample}}\cdot\theta_{\text{out-of-sample}}^{*}\cdot\bar{\epsilon}_{\text{out-of-sample}}-\left(\theta_{\text{in-sample}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in-sample}}^{2}$
Appendix C. Movement Prediction with an Exogenous Variable
A simple but effective method for generating movement prediction involves leveraging the co-movement between the target variable and an exogenous variable. The selection of the exogenous variable is typically guided by domain knowledge, and the movement prediction of the target variable, denoted as $\hat{D}_{t}^{\text{target}}$, can be simply obtained as:
$\hat{D}_{t}^{\text{target}}=\hat{D}_{t}^{\text{exo}},$ (C1)
where $\hat{D}_{t}^{\text{exo}}$ represents the movement prediction of the exogenous variable at time step $t$. In particular, if the value of the exogenous variable is recorded sufficiently ahead of the target variable at the same time step, such as in financial markets operating in different time zones, the actual movement of the exogenous variable can directly serve as the movement prediction for the target variable:
$\hat{D}_{t}^{\text{target}}=D_{t}^{\text{exo}}=\begin{cases}1,&\text{if }y_{t}^{\text{exo}}>y_{t-1}^{\text{exo}};\\ -1,&\text{if }y_{t}^{\text{exo}}\leq y_{t-1}^{\text{exo}};\end{cases},$ (C2)
where $D_{t}^{\text{exo}}$ is the actual movement of the target variable at time step $t$. $y_{t}^{\text{exo}}$ and $y_{t-1}^{\text{exo}}$ denotes the actual values of the exogenous variable at time steps $t$ and t1t-1, respectively. The in-sample movement prediction accuracy of the target variable, denoted as $ACC_{\text{in-sample}}$, can be obtained as follows:
$ACC_{\text{in-sample}}=\frac{1}{n}\sum\limits_{i=1}^{n}{1(D_{i}^{\text{exo}}=D_{i}^{\text{target}})},$ (C3)
where $n$ is the total number of time steps in the in-sample set and $D_{i}^{\text{exo}}$ and $D_{i}^{\text{target}}$ are the $i\text{th}$ actual movements of the exogenous and target variables within the in-sample set, respectively. $1(D_{i}^{\text{exo}}=D_{i}^{\text{target}})$ is an indicator function that equals 1 if the movement prediction of the exogenous variable matches the movement prediction of the target variable and 0 otherwise. Only when the movement prediction accuracy is greater than a predefined threshold (e.g., 0.55) is the movement prediction considered meaningful and suitable for using the MPANF at the second stage.